A projectile is launched at an initial angle ?0 above the horizontal from the edge of the top of a cliff
that is 15 m high. It hits the ground below, at the level of the base of the cliff, 230 m from the cliff,
after a flight time of 4.2 s.
a) What was the initial horizontal component of the projectile’s velocity?
b) What was the initial vertical component of the projectile’s velocity?
c) What were the initial angle ?0 and initial speed ?0?
d) What was the projectile's speed just before it struck the ground?
Answer #1

The vertical velocity would vosinθ  and the horizontal velocity would be vocosθ

A)

Initial horizontal component of velocity

Total horizontal distance travelled

R = 230 m

Time taken to travel this distance

t = 4.2 s

R/t = vocosθ

Vocosθ = 230/4.2

Vocosθ = 54.8 m/s

B)
Initial vertical component of velocity

Total vertical distance travelled

H = ut - ½gt2

15 = (vosinθ x 4.2) - 0.5 x 9.8 x 4.2

Vosinθ =35.58/4.2

vosinθ = 8.47 m/s

C)

Initial angle

Vosinθ/Vocosθ = tanθ

Tanθ = 35.58/230

θ = 8.79o

Initial speed

Vo2 = (vosinθ)2 + (Vocosθ)2

Vo2 = (35.58/4.2)2 + (230/4.2)2

Vo = 55.4 m/s

D)

V = u - gt

V = 8.47 - 9.8 x 4.2

V = -32.7 m/s ( The minus sign just indicates that the final velocity is downwards)